OpenStudy (experimentx):

Let, \( p = {\partial z \over \partial x} \) and \( q = {\partial z \over \partial y}\) Find the solution:- \[ 1) {y^2zp \over x} + xzq = y^2\] \[ 2) (x+2z)p + (4zx-y)q = 2x^2+y\]

OpenStudy (amistre64):

what defines a "solve"?

OpenStudy (experimentx):

find the solution ..

OpenStudy (amistre64):

are these exact equations by chance?

OpenStudy (experimentx):

first order linear PDE ... i know how to solve them. just don't understand the solution.

OpenStudy (experimentx):

according to book ... solution of 1) \( f(x^2-z^2, x^3-y^3) = 0\) 2) \( x^2-y-z =f(xy-z^2)\) I have difficulty interpreting the last step.

OpenStudy (experimentx):

*the solution are

OpenStudy (amistre64):

\[ {y^2z \over x}dx + xz~dy = y^2\] \[ {2y~z \over x} ;~ z\] im having difficulty with the first step, so you one up on me :)

OpenStudy (experimentx):

\[ \huge {y^2z \over x}{\partial z \over \partial x} + xz {\partial z \over \partial y} = y^2\]

OpenStudy (experimentx):

Problem is i don't understand why the solution are of types  u and v are solution of